Presentation is loading. Please wait.

Presentation is loading. Please wait.

Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/2006 1 VIRGO commissioning progress Paolo La Penna for the VIRGO collaboration.

Similar presentations


Presentation on theme: "Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/2006 1 VIRGO commissioning progress Paolo La Penna for the VIRGO collaboration."— Presentation transcript:

1 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ VIRGO commissioning progress Paolo La Penna for the VIRGO collaboration European Gravitational Observatory

2 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Status in December 2005 l Virgo shutdown had started after C7 (19 September 2005) l The new IB had been installed in the IB tower (end November 2005 l The fake dihedron was still mounted on the bench l The Nd:YAG beam (up to 5 W for the moment) had been sent into the bench: some light (with even some interference) was visible at the end of the MC l The optics on the bench were being realigned using the IB actuators: this operation was performed staying inside the tower, with the laminar flux on and the LC closed l The alignment of the IMC was going on (http://www.ego-gw.it/ILIAS-GW/WP1docs/lapenna_ ppt)

3 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Since then: l The IMC was fully aligned in vacuum at about mid December. l IMC locked at end of l Beam at the end of the ITF at begin of January l ITF alignment at mid of January. l First ITF cavities lockings at end of January. l Full RFC alignment at end of February. l Full locking of the ITF going on now (mid of March). The full restart of the ITF, after IB installation, took about three months

4 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ I – Injection system restart

5 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Operation performed on the Injection System l Alignment and locking of the IMC in vacuum l Alignment of the IB optics in vacuum l Alignment of the beam on the ITF and telescopes tuning l Alignment and locking of the RFC l Setup and alignment of the EIB l BMS l IMC matching improvement l IMC automatic alignment l Power stabilization l RFC automatic alignment

6 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Changes in Virgo after September shutdown IB+RFC EIB PR mirror

7 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Injection system Lower stage Upper stage

8 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Difference old-new IB Old IBNew IB

9 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC alignment: december 2005 HeNe autocollimation on the MC mirror Acting on the IB alignment (checked with a ccd)

10 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC alignment: december 2005 YAG beam direct beam centering on the MC mirror acting on the external steering mirrors laser

11 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC alignment: december 2005 MC mirror alignment until IMC resonances are visible laser

12 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC locking: december 2005 IMC locking Feedback on laser pzt using the reflection as a trigger laser P out 8 W

13 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ ITF alignment: 13 january 2006

14 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC alignment and locking: february 2006 l The Reference Cavity alignment has taken more time than expected l A first attempt to align it using the 20 W beam failed l A second attempt after having aligned a probe beam coming from the opposite direction, and looking at the RFC transmission, led to a poor RFC alignment (but it allowed to lock and start working on the ITF) l A third attempt overlapping the probe beam and the 20 W beam was successful l From the beginning to the end almost one month was necessary

15 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ YAG from inverse path: fringes Secondary YAG beam 1 25 Hz pzt scanning lProbe beam from the inverse path lProbe beam superposed to a HeNe passing through the RFC in the lower part lCamera looking at the coils frame (hit by the RFC transmission): mode flashing visible ccd photodiode ccd lCCD and PhD looking at RFC reflection

16 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Contrast improvement and alignment YAG probe Reflected beam Contrast = 0.7

17 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC alignment using PR alignment

18 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC alignment using PR alignment

19 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Near Field of M13 and Far Field 1m-f FF NF Screens

20 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ M14 movement 1m-f NF Move M14 and observe overlapping with NF ccd FF

21 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ M13 movement 1m-f Move M14 and observe overlapping with NF ccd Correct with M13 to overlap on the FF ccd NF FF

22 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ After alignment: the 20 W beam resonates in the RFC ccd PhD

23 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ The RFC transmission is not visible with the RFC aligned lInverse path beam and direct beam are being superposed below acting on M13 and M4 mirrors lWhen superimposed both beams should resonate ccd PhD

24 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC locked, reflected 20 W YAG beam: 1 March 2006 Contrast = 54 % (it was 80% before the shutdown) Good alignment + longitudinal mismatching

25 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC alignment: conclusion l The RFC alignment was more difficult than expected: the cavity axis was not where it was expected, maybe the RFC was not aligned inside its tank, or it had moved slightly during new IB assembling l The RFC is locked in a stable way, with the beam transmitted by the IMC: more stable with the alignment with respect to the previous setup l We cannot see the RFC transmission

26 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Further activities on RFC l RFC locking filter change: a new filter has been implemented, which allows a reduction of the frequency noise of the prestabilized beam l RFC automatic alignment: this activity is going on; it will allow to keep the beam aligned on the RFC using the two closed loop pzt placed before the RFC. This allows shifts of the beam on the upper part, in order to better align the upper part optics, without losing the RFC alignment

27 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Problems encountered DPS misaligned PM stuck TS damaged Wrong PSD electronics Beam clipped RFC_trans clipped Several problems/failure have occurred: it has been necessary to open the IB tower twice

28 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Further activities involving the injection system l Beam Steering System l IMC matching improvement l IMC automatic alignment restarted l Input power stabilization (the photodiode measuring the IMC transmission has been changed) l RFC automatic alignment l PSD on the IB working (new homemade electronics) After the restart of the IB, some new features and other activities have been performed:

29 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ BMS: DC quadrants, ABP actuation ABPs (on laser bench) NF QF FF QF l BMS allows to keep beam aligned with respect to DC_quadrants placed on the External Injection Bench. l The signal from the DC quadrants is sent to the ABP (Automatic Beam Positioning) mirrors (pzt steered mirrors). l It allows a stable and repetitive positioning of the beam entering into the IMC; l It allows performing pure translations and angle of the beam.

30 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC matching l With the help of BMS, IMC matching has been improved: 96% matching (it was 82%), 91% contrast (should correspond to about 1500 ppm losses in the IMC). l IMC transmission is now about 8 W (at about the same level as before the shutdown) l This power is entering into the ITF (we had 0.7 W entering into the ITF before the shutdown)

31 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IMC automatic alignment l The AA of the MC is working again as before the shutdown (for the moment acting only on the MC mirror, correcting the FF_vert and the NF_horiz degrees of freedom. l The (new) AA of the IB is ready to be implemented: it should replace the LC of the IB

32 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Power Stabilization Data from spectrum analyser is 50 times lower in the 100Hz region than dataDisplay measurent (above) (to be investigated) Master laser Diode current IMC_trans HF

33 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ RFC automatic alignment l Closed on the vertical DOF (1 Hz bandwidth) l Horizontal DOF not working (M16_PZT horiz. not working properly) M16 M15 Wavefront sensing quadrants on the EIB RFC AA now separated from IMC AA

34 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ New PR mirror: monolithic and higher R (0.95 vs 0.925) Comparison between the old (black) and new monolithic (red) PR mirror

35 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ II- ITF restart

36 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IB LC after RFC locking The IB LC filters were changed after restart: frequency noise was reintroduced

37 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Frequency noise after RFC locking (ITF misaligned) IB transfer function Frequency noise

38 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ IB LC filter upgrade IB in high noise controlIB in low noise control Blue curves: MC in high/low noise (factor 100 G difference) Red curves: MC in high/low noise (factor 100 G difference)

39 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ MC-RFC locking filter: after restart Too small G SSFS couldnt engage

40 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Frequency noise after RFC locking (ITF misaligned) Purple: C7 Black: after new IB RFC locking

41 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ New MC-RFC locking filter upgrade

42 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ New frequency noise: SSFS can engage

43 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Frequency noise: C7 (beam attenuated) vs now C7 Now

44 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ ITF restart: locking l Photodiodes gains readjusted for the higher power (10 times more than before shutdown) l Automatic alignment restarted: Tested the drift control of the two cavities Quadrant photodiodes gains adjusted Not gone further because it needs the full locking of the ITF l PR-NI l PR-NE l CITF l NE-WE-BS l NA l WA l Recombined l SSFS l Variable Finesse with PR misaligned l …. Now the full recycled ITF is being locked Since end of January:

45 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ SSFS L1L1 L2L2 l2l2 l1l1 l0l0 B1p_DC B5_ACp SSFS using CARM for the Beam Frequency stabilization B2_3f_ACp PR (PRCL) B1p_DC BS (MICH) B5_ACp SSFS (CARM) B8_ACp NE-WE (DARM) B8p B2_3f_ACp

46 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ SSFS restart The restart of SSFS required some gain tuning because: l RFC locking filter was different l Frequency noise was different l RFC gains are different After tuning readjustment, IB LC filters and RFC locking filter SSFS is now working properly

47 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Towards the dark fringe L1L1 L2L2 l2l2 l1l1 l0l0 STEP 7 B5_ACp SSFS using CARM DF 0.05 B8p B2_3f_ACp B1p_DC B2_3f_ACp PR (PRCL) B1p_DC BS (MICH) B5_ACp SSFS (CARM) B8_ACp NE-WE (DARM)

48 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ ITF locking l Status on 16th March: ITF locked up to step 7 (0.05 of the dark fringe) Here we are

49 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Parabolic telescope T2: M6 mirror misalignment 3 closed loop picomotors Tilted: tx (650 rad) and ty (850 rad)

50 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Effect of M6 misalignment M6 aligned (beam at 3 km) M6 misaligned (beam at 3 km) ( x=650 rad y=850 rad)

51 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Resulting beam shape Beam size at NE: wx= 60mm wy= 60mm Beam size on Cam5: wx= 0.7mm wy= 0.42mm To fit these measurements, the beam parameters inside ITF with respect to BS position are (perfect matching for w=21.3, z=0): wx= 15mm, zx= -400m wy= 23mm, zy= 800m The matching on the FP cavities should be around 89% in these conditions: confirmed by measurement (88%). Beam parameters before shutdown: wx= 18mm, zx= -500m wy= 18mm, zy= 500m)

52 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Short telescope T1 centering Y L1 f=130 L2 f=-80 D=60mm ccd

53 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Beam shape after the Faraday isolator y=4 mm w/w = 9% y=2 mm w/w = 2%

54 Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/ Conclusion l The restart of the Injection System has required more than two months l The new Injection System is now working l The ITF is close to final locking with full power (8 W) Some problems involving the new IB remain open: l The RFC matching is poor and the RFC transmission is not sent out of the IB tower l The astigmatism of the beam has to be corrected For the moment the effort is in achieve the full locking again and, since we can go on in these conditions, to pospone the solving of this problems to a later moment


Download ppt "Paolo La Penna Virgo Commissioning StatusN5-WP1 8 TH MEETING, Frascati, 20/03/2006 1 VIRGO commissioning progress Paolo La Penna for the VIRGO collaboration."

Similar presentations


Ads by Google